OpenStudy (anonymous):

Find the indicated dot product. r = <9, -7, -8>, v = <3, 4, 7>, w = <6, -9, 7> v ⋅ w <-27, 28, 56> <18, -36, 49> 18 31

OpenStudy (anonymous):

At first i thought the formula i needed to use was \[v*w= (v _{1}*w _{1})+(v _{2}*w _{2})+(v _{3}*w _{3}) \] But I'm not getting the right answer.. i figure im doing something wrong

OpenStudy (bibby):

that's the cross product

OpenStudy (anonymous):

oh.. lol that explains alot

OpenStudy (bibby):

you can use \(\large \overrightarrow u \cdot \overrightarrow v = u_x*v_x+u_y*v_y\)

OpenStudy (bibby):

as usual, wikipedia does it better "his operation can be defined either algebraically or geometrically. Algebraically, it is the sum of the products of the corresponding entries of the two sequences of numbers."

OpenStudy (anonymous):

omg thank you so much !!

OpenStudy (bibby):

just wondering what the answer is, asking for a friend

OpenStudy (anonymous):

umm gimme a sec. i had already submitted the exam and got it wrong .. ill do the math though

OpenStudy (bibby):

it's cool if you don't want to, I just wanted to make sure you understood the concept

OpenStudy (anonymous):

yeah , no i dont mind.. i still dont understand though.. theres 3 numbers though ? what do i do with the 3rd? the formula only lets me do 2 number

OpenStudy (bibby):

it's the same thing but x,y,z instead of x y notice the wikipedia articles wording. it says componentwise there is no formula, that's just the formula for the case where u and v have 2 components

OpenStudy (anonymous):

ooohhhhh alright .. gimme a sec again lol

OpenStudy (bibby):

oh, same thing for 3 vectors btw

OpenStudy (unklerhaukus):

\[v = \langle v_0, v_1, v_2\rangle,\quad w = \langle w_0, w_1, w_2\rangle \] \[v\cdot w=\langle v_0, v_1, v_2\rangle\cdot \langle w_0, w_1, w_2\rangle\\ \qquad=v_0w_0+v_1w_1+v_2w_2\]

OpenStudy (anonymous):

i got 31

OpenStudy (bibby):

so \(u*v*w = u_x v_x w_x + u_y v_yw_y+u_zv_zw_z\)

OpenStudy (unklerhaukus):

the dot product of two vectors, is simply the sum of the products of the corresponding componets

OpenStudy (unklerhaukus):

\[\checkmark\]

OpenStudy (anonymous):

Thank you both of you ! Big help !